A 275N crate is sliding down a 35 degrees incline. If the force of friction along the incline is 96N, what is the acceleration of the crate?


1
Expert's answer
2021-03-14T19:20:33-0400

The projection of the gravity force on the crate along the incline (the force that pushes the crate down) is:


"F_{down} = 275\cdot \sin 35\degree"

The force that pushes it up along the incline is the friction force:


"F_{up} = 96N"

Thus, according to the second Newton's law, their difference is:


"F_{down} - F_{up} = ma\\
a = \dfrac{F_{down} - F_{up} }{m}"

where "m" is the mass of the crate, and "a" is its acceleration.

The mass can be found from the weight:


"m = \dfrac{275N}{9.8N/kg}"

where "9.8N\/kg" is the gravitational acceleration.

Thus, obtain:


"a = \dfrac{ 275\cdot \sin 35\degree}{\dfrac{275N}{9.8N/kg}} - \dfrac{96N}{\dfrac{275N}{9.8N/kg}} \approx 2.2m/s^2"

Answer. 2.2 m/s^2.